The rate constant for a reaction of zero order in A is 0.0030 mol L-1 s -1. How long will it take for the initial concentration of A to fall from 0.10 M to 0.075 M? 

Grade:12

2 Answers

Nicho priyatham
8 years ago
since first order (af-a)/t =k
so t= (af-a)/k=25/3 seconds
plz approve if useful
grenade
8 years ago
for a 0 order reaction
k= [R0]-[R]/t
the then on substitutig the values we will get the answer as 25/3 = 8.33